# Adaptation to extreme stress under the growth-survival fitness trade-off

### v1

@@ Abstract
Microbial adaptation to extreme stress, such as starvation, antimicrobial exposure, or freezing often reveals fundamental trade-offs between survival and proliferation. Understanding how populations navigate these trade-offs in fluctuating environments remains a central challenge. We develop a quantitative model to investigate the adaptation of populations of yeast (Saccharomyces cerevisiae) subjected to cycles of growth and extreme freeze-thaw stress, focusing on the role of quiescence as a mediator of survival. Our model links key life-history traits—growth rate, lag time, quiescence probability, and stress survival—to a single underlying phenotype, motivated by the role of intracellular trehalose in the adaptation of yeast to freeze-thaw stress. Through stochastic population simulations and analytical calculation of the long-term growth rate, we identify the evolutionary attractors of the system. We find that the strength of the growth-survival trade-off depends critically on environmental parameters, such as the duration of the growth phase. Crucially, our analysis reveals that populations optimized for growth-stress cycles can often maintain viability alongside growth-optimized populations even in the absence of stress. This demonstrates that underlying physiological trade-offs do not necessarily translate into fitness trade-offs at the population level, providing general insights into the complex interplay between environmental fluctuations, physiological constraints, and evolutionary dynamics.
1 Introduction
@@
Adaptation to extreme stress in microbial systems often occurs at a cost to their ability to grow quickly under ambient conditions [1, 2]. Specifically, an increase in survival under oxidative stress, glucose limitation, freeze-thaw, and heat exposure redirects resources from growth pathways to those dedicated to damage reduction and repair [3]. While trade-offs are not universal across all environmental niches [4, 5], they are a common outcome of adaptation to extreme stressors. These evolutionary trade-offs directly influence the emergence of generalists or specialists depending upon the extremity and frequency of environmental change [6, 7, 8].
@@
Here we present a model of adaptation of Saccharomyces cerevisiae to repeated cycles of growth and nutrient exhaustion followed by freeze-thaw stress, based upon experiments covered in previous work [9]. Evolution under these conditions produces a clear fitness trade-off: freeze-thaw survival increases while the growth rate in ambient conditions decreases. Our model simulates the evolution of a yeast population through these repeated cycles and explicitly addresses the roles of quiescence and lag time.
@@
Further, when yeast is subject to freeze-thaw stress without allowing the population to reach stationary phase, the population crashes, suggesting that quiescence acts as a helpful step to the evolution of freeze-thaw tolerance. Recent work in yeast has shown that stress response often follows a general strategy rather than evolving to specific stress conditions [1, 2, 20]. For instance, yeast evolved under aerobic glucose limitation show few trade-offs with growth in other carbon-limited environments [2]. While the fitness of the evolved strain increases in comparison to yeast growing in rich media, these fitness improvements can arise from multiple mutational paths. This general stress response is most evident at the level of the phenotype and includes producing higher levels of stationary phase trehalose, thickening of the cell wall, a reduction in size [9], and is underpinned by large-scale, genome-wide transcriptional reprogramming [20, 21]. Therefore, while our model is grounded in a specific experimental system, its results can generalize to any microbial population evolving under cycles of growth and stress exposure.
@@
Our model highlights the role of quiescence in adapting to extreme stress. Yeast populations undergo a diauxic shift following glucose depletion, where cells reduce their growth rate to readjust their metabolism for the subsequent phase of slow, respiratory growth. After nonfermentable carbon sources have been consumed, the population enters stationary phase. Stationary phase cultures are complex and heterogeneous, composed of a large fraction of quiescent, long-lived, largely daughter and young mother cells, and a nonquiescent subpopulation of cells [22, 23, 24, 25, 17]. The non-quiescent subpopulation eventually perishes under conditions of long term starvation. Quiescence is a distinct, growth-arrested phase of the cell cycle [22, 26, 27].
@@
Thus, quiescence of yeast cells in conditions of starvation is a milder stress response which could prime them for harsher conditions, enabling survival during long periods of starvation [28]. Quiescent cells share several key characteristics with response to extreme stress, including smaller, denser cells with thicker cell walls [29, 25], and also show increased resistance to diverse environmental stresses [28, 30, 31]. However, the exact relationship between quiescence – a separate cell cycle phase involving stochastic switching – and canonical cellular stress response remains unclear [30, 23]. Distinct from stress-resistant cells, quiescent cells show growth arrest and stochastic switching between the growth cycle and quiescence. The dynamics of entry into and exit from quiescence has been shown to be a noisy process that couples deterministic memory effects of preceding cell cycle with stochastic switching [32]. Yet, the large scale rearrangement of gene expression, metabolic state and cellular processes that takes place in quiescence may show similarity to stress response in yeast [33, 3, 34, 35].
@@
In Section 2 we introduce our model of microbial evolution under growth-starvation-stress cycles that treats the switch to quiescence as a stepping stone to stress-resistance. The model qualitatively captures the evolution of yeast populations under cycles of growth, starvation, and freeze-thaw stress. We specifically confirm its predictions for the evolutionary trajectory of the central phenotype, the amount of intracellular trehalose in stationary phase, with experimental results. In Section 3 we calculate the long term growth rate of a population under growth-stress cycles and use it to calculate the optimum phenotype for different evolution programs. We find that despite the growth-survival trade-off, populations evolved under growth-stress cycles can coexist with those evolved in the absence of stress.
2 Trehalose Mediated Quiescence, Lag Time, Growth and Survival
@@
Entry into quiescence is a stochastic process, while the efficiency of a cell’s exit from quiescence is expressed in the duration of the lag phase upon re-entry into rich media. The duration of the lag phase is also determined by the amount of intracellular trehalose produced in stationary phase, as trehalose is preferentially metabolized to re-enter the cell cycle.
We assume that a single phenotype is able to capture different aspects of a cell’s fitness. We take the primary phenotype to be the amount of intracellular trehalose in the cell when the population enters stationary phase. It is well known that higher trehalose levels correspond to lower cellular damage under physical stress, as well as higher survival fractions under freezing and desiccation. The cellular mechanisms that produce trehalose as part of a broader and general stress response show a trade off with the pathways that promote cell division. Thus, as trehalose production under stressful conditions increases, growth rate of the cell upon reentry into the growth phase decreases. All of this is also observed in our experimental results, and we consider that the stationary phase trehalose levels in the cells can determine the cell’s growth rate, quiescence probability, survival probability, and lag time upon reentry into the cell cycle.
2.1 Methods
We start with a population of $N$ individual cells whose phenotype, given by $\phi$, is drawn initially from a Gaussian distribution with mean $\phi_{0}$ and a narrow peak with standard deviation $\sigma_{\phi}^{0}$. We assume the dependence of each cell’s growth rate (or fecundity per unit time in the simulation) on stationary phase trehalose concentration to be quadratic,
@@ 2.1 Methods
where $g_{0}$ is the highest achievable growth rate and $\tau_{g}$ the amount of trehalose that inhibits growth to effectively zero. Note that our population does not typically reach $\phi$ values close to $\tau_{g}$ during the course of the simulation, and we set $g(\phi)=0$ for $\phi>\tau_{g}$. At each time step in the growth process, we calculate the number of offspring for each cell by applying a ceiling function to $g(\phi)$ calculated according to the above equation, and add them to the population. We consider here a quadratic function for $g(\phi)$ so as to allow the growth rate to fall faster with increasing $\phi$ values at lower values and slower as $\phi$ reaches higher values. Each offspring is also subject to a small mutation, which acts to change the trehalose concentration of the offspring with respect to its parent. This mutational effect on offspring phenotype is drawn from a Pearson Distribution at a mutation rate $m$ for each offspring. The Pearson Distribution is a continuous probability distribution with a long tail defined by its kurtosis. Our distribution has a mean at zero, mutational standard deviation $\sigma_{m}$, a kurtosis $k_{m}$ and no skew. A kurtosis $k=3$ is a Normal distribution, with outliers becoming more probable with increasing kurtosis. For our simulations we use $m=0.8$ and $k_{m}=5$. Most mutations have no effect upon the offspring’s phenotype, preserving the value inherited from the parent cell. Simultaneously, some large effect mutations, both positive and negative are possible.
For each division, we assume a constant amount of a finite resource, $r_{d}$, is consumed, starting with $r_{0}$ resource units. Cells also consume a smaller amount of the resource, $r_{l}$ while they are non-growing in lag phase. The growth process continues till the resource runs out and the population faces resource exhaustion. For our simulations we use $r_{d}=0.5,r_{l}=0.1$ and $r_{0}=10^{5}$ Upon starvation, yeast forms a heterogeneous population of quiescent and non-quiescent cells. We model the switch to quiescence as a stochastic process, where the probability of each cell going into quiescence is given by
@@
where $\tau_{q}$ is the trehalose concentration at which probability of quiescence is $50\%$ and $\gamma_{q}$ is a parameter quantifying how quickly the probability rises to $1$ as $\phi$ increases. The functional form, as well as the $\tau_{q}$ values we chose allows for probability of quiescence to gradually increase with $\phi$ and saturate at a high value. For each cell we decide the fate of the cell by comparing this probability with a randomly generated number. If the probability given by Equation 2 is higher than the random number the cell goes into a quiescent state.
Note that the phenotype of each cell changes from mutations during the growth process. The phenotype corresponds to trehalose concentrations produced by the cell eventually in the stationary phase, which also determines its quiescence probability. Thus mutations accrued in growth process can bias a cell towards or away from quiescence, although no trehalose is produced in the growth phase.
Entry upon quiescence is followed by exposure to extreme stress, which in our experiments is freeze-thaw. Upon exposure to extreme stress, we assume that the non-quiescent population entirely perishes. Among the quiescent cells, the cells with higher trehalose concentration have a higher probability of survival. We model the probability of survival,
@@
where $\tau_{s}$ is the trehalose concentration at which probability of survival under freeze thaw is $50\%$ and $\gamma_{s}$ is a parameter quantifying how quickly the probability rises to $1$ as $\phi$ increases. Again, we choose $\gamma_{s}$ and $\tau_{s}$ such that the probability of survival is close to zero for low values of $\phi$ and then increases relatively quickly in an interval given by $\gamma_{s}$, representing a switch-like transition to having enough trehalose production in each cell to lend the cell resistance to freeze-thaw. In general we expect that $\gamma_{s}<\gamma_{q}$ and $\tau_{s}>\tau_{q}$ since the stress is harsher to cells than nutrient scarcity in stationary phase. Thus, both quiescence and the freeze thaw process shift the mean phenotype of the population to higher values since cells with lower phenotype amounts perish. The functions for growth rate, quiescence probability and survival probability are shown in Figure 1 (a).
When the surviving population is allowed to thaw and re-introduced into a growth medium, the cells with higher trehalose concentrations have a shorter lag time, and begin to grow earlier as trehalose is crucial for re-entry into the cell cycle [16]. Longer lag times mean less resource available for growth because a small amount of resource is consumed even in the lag phase and the total amount consumed before growth depends on the length of the lag phase. We assume the simplest possible functional form, a linear relationship between the lag time and trehalose amounts,
@@
where $l_{0}$ represents the maximum lag time possible, and $\tau_{l}$ the phenotype value at which the lag time reduces to zero. We set $l(\phi)=0$ for all $\phi>\tau_{l}$ for the sake of our simulation, although $\tau_{l}$ is set large enough that $\phi$ does not exceed it in the simulations. In each growth cycle, a phenotype value thus means a shorter lag phase, and more time available for growth, while simultaneously causing a lower growth rate. Lastly, we model the dilution process of the experimental setup by randomly selecting a fraction $d$ of the population. This can introduce additional stochasticity into the evolution process.
2.2 Results
@@
When the population starts growth after exit from lag phase with lower values of mean trehalose, the effect of a decreasing growth rate is not sufficient to stop the increase in mean trehalose of the population caused by a shorter lag phase. In the growth phase, cells with higher trehalose are able to produce a larger number of offspring and take over the population.
@@
As the population undergoes a complete round of growth and nutrient exhaustion, we have a resultant heterogeneous population, with cells with higher $\phi$ having a higher probability of entering quiescence. Thus, the distribution of quiescent cells now has a higher mean amount of trehalose production than the entire population taken together. A fraction of theses quiescent cells are able to survive freeze-thaw stress. The cells that are now able to survive are now a larger proportion of the entire population because of this upward shift in $\phi$. Hence, our simulation includes both a component of trehalose-dependent stochastic switching to quiescence, as well as growth, lag and survival mediated selection for trehalose levels. The trade-off between the decrease in growth rate and the positive effect on lag time and survival from increasing trehalose sets the final equilibrium value of trehalose in the simulations. The effect of a lower growth rate begins to dominate at larger trehalose values, causing the mean trehalose of the population to stagnate.
@@
We see an increase in the phenotype and survival as the evolution process progresses, before settling at an equilibrium value. This is shown in Figure 1 (b). The blue curves represent individual replicate populations, while the red curve represents the average over $100$ replicates. The behavior of populations is highly stochastic in the early stages of the evolutionary process, due to small numbers of cells that survive freeze-thaw stress, but the equilibrium value they settle to remains conserved across all replicates.
@@
Figure 2 (a) and (b) show the evolution of the populations mean phenotype and the resultant survival fractions over time. We compare them to the trehalose evolution curve (c) and the survival fraction curve (d) obtained from the experimental evolution of yeast under freeze thaw. Both sets of curves show qualitative agreement. The rapid increase in the survival rate is caused by the large tailed mutational distribution we use which biases evolution towards large jumps. We note also the delay in increase of both the survival curves, which arises, in the case of our simulation, from the sigmoidal dependence of the survival probability upon trehalose concentration.
@@
We investigate the effect of increasing $g_{0}$ on equilibrium values of the phenotype and survival fraction in Figure 3 (a) and (b). Increasing $g_{0}$ increases the number of offspring at each time point, leading to the resource being exhausted faster, thus less time is spent by the population in its growth phase. This causes the effect of selection from growth to decrease, allowing the phenotype to increase to higher values and biasing the population towards maximizing survival. Thus, we see an increase in both the equilibrium value of the phenotype as well as survival fraction as $g_{0}$ increases.
@@
Hence we find that the amount of time spent in the growth phase, the maximum growth rate, among other parameters can affect the equilibrium value of the phenotype and survival by changing the strength of the trade-off. We write down the exact form of this dependence in Section 3 by estimating the long term growth rate of the population.
3 Long Term Growth Rate and Comparing Fitness
We calculate the long term growth rate of the evolving population to help us to analytically recover some of the results of our model. Consider a population of cells evolving under growth-stress cycles given by the distribution $n(\phi,t)$. As in the last section, the growth rate, probability of survival under stress and lag time all depend on the phenotype of each cell. Through the growth phase, the number of cells with phenotype $\phi$ at time $(t+1)$ is given by $g(\phi)n(\phi,t)$, where $g(\phi)$ is the growth rate at phenotype $\phi$. If the population is allowed to grow for $T_{g}$ time till the resource runs out, at the end of the growth phase the number of cells at each phenotype is

### v2

@@ Abstract
Microbial adaptation to extreme stress, such as starvation, antimicrobial exposure, or freezing often reveals fundamental trade-offs between survival and proliferation. Understanding how populations navigate these trade-offs in fluctuating environments remains a central challenge. We develop a quantitative model to investigate the adaptation of populations of yeast (Saccharomyces cerevisiae) subjected to cycles of growth and extreme freeze-thaw stress, focusing on the role of quiescence as a mediator of survival. Our model links key life-history traits—growth rate, lag time, quiescence probability, and stress survival—to a single underlying phenotype, motivated by the role of intracellular trehalose in the adaptation of yeast to freeze-thaw stress. Through stochastic population simulations and analytical calculation of the long-term growth rate, we identify the evolutionary attractors of the system. We find that the strength of the growth-survival trade-off depends critically on environmental parameters, such as the duration of the growth phase. Crucially, our analysis reveals that populations optimized for growth-stress cycles can maintain viability alongside growth-optimized populations even in the absence of stress. This demonstrates that underlying physiological trade-offs do not necessarily translate into fitness trade-offs at the population level, providing general insights into the complex interplay between environmental fluctuations, physiological constraints, and evolutionary dynamics.
preprint: APS/123-QED
I Introduction
@@
Adaptation to extreme stress in microbial systems often occurs at a cost to their ability to grow quickly under ambient conditions [1, 2]. Specifically, an increase in survival under oxidative stress, glucose limitation, freeze-thaw, and heat exposure redirects resources from growth pathways to those dedicated to damage reduction and repair [3]. While trade-offs are not universal across all environmental niches [4, 5], they are a common outcome of adaptation to extreme stressors. These evolutionary trade-offs directly influence the emergence of generalists or specialists depending upon the extremity and frequency of environmental change [6, 7, 8]. Fitness trade-offs can manifest in distinct life history traits by affecting growth rates, cellular response to starvation, survival and characteristics of the lag phase in distinct ways. This leads to a complex interplay between environmental variation, population dynamics and underlying physiological constraints during the process of evolution. The switch to quiescence is a crucial part of the population response to starvation, and its role in evolutionary adaptation remains relatively unexplored, perhaps with an exception in the case of bacterial persistor cells [62].
@@
Here we present a model of adaptation of Saccharomyces cerevisiae to repeated cycles of growth and nutrient exhaustion followed by freeze-thaw stress, based upon experiments covered in previous work [9]. Evolution under these conditions produces a clear fitness trade-off: freeze-thaw survival increases while the growth rate in ambient conditions decreases, and additionally the duration of lag time decreases. Our model simulates the evolution of a yeast population through these repeated cycles and explicitly addresses the roles of quiescence and lag time in the adaptation process.
@@
Further, when yeast is subject to cycles of growth and freeze-thaw stress without allowing the population to reach stationary phase, the population crashes, suggesting that quiescence acts as an important step to the evolution of freeze-thaw tolerance. Recent work in yeast has shown that stress response often follows a general strategy rather than evolving to specific stress conditions [1, 2, 20]. For instance, yeast evolved under aerobic glucose limitation shows few trade-offs with growth in other carbon-limited environments [2]. While the fitness of the evolved strain increases in comparison to yeast growing in rich media, these fitness improvements can arise from multiple mutational paths. This general stress response is most evident at the level of the phenotype and includes producing higher levels of stationary phase trehalose, thickening of the cell wall, a reduction in size [9], and is underpinned by large-scale, genome-wide transcriptional reprogramming [20, 21]. Therefore, while our model is grounded in a specific experimental evolution system, its results can be generalized to different microbial populations evolving under cycles of growth and stress exposure.
@@
We highlight the role of quiescence in adapting to extreme stress. Yeast populations undergo a diauxic shift following glucose depletion, where cells reduce their growth rate to readjust their metabolism for the subsequent phase of slow, respiratory growth. After nonfermentable carbon sources have been consumed, the population enters stationary phase. Stationary phase cultures are complex and heterogeneous, composed of a large fraction of quiescent, long-lived, largely daughter and young mother cells, and a nonquiescent subpopulation of cells [22, 23, 24, 25, 17]. The non-quiescent subpopulation eventually perishes under conditions of long term starvation. Quiescence is a reversible exit from the cell cycle in the stationary phase, distinct from growth arrest due to nutrient starvation [22, 26, 27].
@@
Thus, quiescence of yeast cells in conditions of starvation is a milder stress response which could prime them for harsher conditions, enabling survival during long periods of starvation [28]. Quiescent cells share several key characteristics with response to extreme stress, including smaller, denser cells with thicker cell walls [29, 25], and also show increased resistance to diverse environmental stresses [28, 30, 31]. However, the exact relationship between quiescence and canonical cellular stress response remains unclear [30, 23]. Distinct from stress-resistant cells, quiescent cells show cell division arrest and stochastic switching between the cell cycle and quiescence. The dynamics of entry into and exit from quiescence has been shown to be a noisy process that couples deterministic memory effects of preceding cell cycle with stochastic switching [32]. Yet, the large scale rearrangement of gene expression, metabolic state and cellular processes that takes place in quiescence may show similarity to stress response in yeast [33, 3, 34, 35].
@@
In Section II we introduce our model of microbial evolution under growth-starvation-stress cycles that treats the switch to quiescence as a stepping stone to stress-resistance. The model qualitatively captures the evolution of yeast populations under cycles of growth, starvation, and freeze-thaw stress. We specifically confirm its predictions for the evolutionary trajectory of the central phenotype, the amount of intracellular trehalose in stationary phase, with experimental results. Using analytic calculations, we estimate the long term growth rate of a population under growth-stress cycles in Section III and use it to calculate the optimum phenotype for different evolution programs. We find that despite the growth-survival trade-off, populations evolved under growth-stress cycles can have comparable fitness to those those evolved in the absence of stress. This suggests the possibility of coexistence. We explore this suggestion using simulations of finite resource competitive growth of the two optimal strains.
II Trehalose Mediated Quiescence, Lag Time, Growth and Survival
Figure 1: Schematic showing the first round of the evolution simulation. (a) An initial population drawn from a Gaussian distribution in $\phi$ is selected and allowed to undergo resource limited growth till it reaches a stationary phase at resource exhaustion. (b) The stationary phase population distribution is broader than the initial distribution due to mutations, and its mean is shifted towards the left due to higher growth rates at lower $\phi$ values. (c) This population then differentiates into quiescent and non-quiescent cells, with the probability of quiescence imposing a lower threshold in $\phi$ on the population distribution. (d) The population is subsequently exposed to extreme stress and a smaller fraction of the quiescent population survives above a second threshold value of $\phi$ imposed by the survival probability. (e) As the population is given a resource source again, cells in the distribution exit lag phase and begin growth. Higher $\phi$ values mean lower lag times, but also lower growth rates. The next stationary phase population is a result of these two competing effects. The simulation then continues with the population again differentiating into quiescent and non-quiescent cells.
@@ II Trehalose Mediated Quiescence, Lag Time, Growth and Survival
Entry into quiescence is a stochastic process, while the efficiency of a cell’s exit from quiescence is expressed in the duration of the lag phase upon re-entry into rich media. The duration of the lag phase is also determined by the amount of intracellular trehalose produced in stationary phase, as trehalose is preferentially metabolized to re-enter the cell cycle. Further, higher trehalose levels correspond to lower cellular damage under physical stress, as well as higher survival fractions under freezing and desiccation. Lastly, the cellular mechanisms that produce trehalose as part of a broader and general stress response show a trade off with the pathways that promote cell division. Thus, as trehalose production under stressful conditions increases, growth rate of the cell upon reentry into the growth phase decreases. All of this is observed in our experimental results, and we postulate that the stationary phase trehalose levels in a cell can determine the cell’s growth rate in exponential phase, quiescence probability, survival probability, and lag time upon reentry into the cell cycle.
For building our model and simulations, we assume that a single phenotype, $\phi$ is able to capture different aspects of a cell’s fitness, including growth rate, probability of making a switch into quiescence in stationary phase, survival probability in extreme stress and the lag time of the cell once reintroduced into fresh media. This primary phenotype corresponds to the amount of intracellular trehalose in the cell when the population enters stationary phase.
We start with a discrete population of NN individual cells whose phenotype, given by $\phi$, is drawn initially from a continuous Gaussian distribution with mean $\phi_{0}$ and a narrow peak with standard deviation $\sigma_{\phi}^{0}$. Note that each cell has a phenotype given by $\phi$, and the population is characterized by a distribution in $\phi$. This is shown in Figure 1(a).
Our simulation is divided into four phases; (1) growth phase, (2) quiescence phase, (3) survival phase, and (4) lag phase. To mimic the experimental setup, the first phase in our simulation is a growth phase where the population is introduced into fixed resource, $r_{0}$. Each time step in this phase of the simulation is a new generation of the population, and we assume non-overlapping generations. Each cell in the population has a different growth rate determined by its phenotype. We assume the dependence of growth rate (or fecundity per unit time in the simulation) on $\phi$ to be quadratic,
@@
where $g_{0}$ is the highest achievable growth rate and $\tau_{g}$ the amount of trehalose that inhibits growth to effectively zero. Note that although our population does not typically reach $\phi$ values close to $\tau_{g}$ during the course of the simulation, we set $g(\phi)=0$ for $\phi>\tau_{g}$.
At each generation time step in the growth process, we calculate the number of offspring for each cell using $g(\phi)$. We use a quadratic function for $g(\phi)$ to allow the growth rate to fall faster with increasing $\phi$ at lower values and slower as $\phi$ reaches higher values. Since the number of cells is discrete while the growth function and phenotype are continuous functions of $\phi$, we apply a ceiling function to $g(\phi)$ to calculate the number of offspring for each cell add them to the next population generation.
Each offspring is subject to a small mutation, which acts to change the phenotype, $\phi$ of the offspring with respect to its parent. This mutational effect on offspring phenotype is drawn from a Pearson Distribution at a mutation rate $m$ for each offspring. The Pearson Distribution is a continuous probability distribution with a long tail defined by its kurtosis. Our distribution has a mean at zero, mutational standard deviation $\sigma_{m}$, a kurtosis $k_{m}$ and no skew. A kurtosis $k=3$ is a Normal distribution, with outliers becoming more probable with increasing kurtosis. For our simulations we use $m=0.8$ and $k_{m}=5$. Since the distribution of mutational effects has no skew, the median of the distribution is at 0. This means that most mutations have no effect upon the offspring’s phenotype, preserving the value inherited from the parent cell. Simultaneously, because of our choice of a high kurtosis, some large effect mutations, both positive and negative are possible. Although the mutations almost never drive $\phi$ to be a negative value, we explicitly reject mutations that cause $\phi<0$ in our simulations.
For each cell division event, we assume a constant amount of the finite resource, $r_{d}$, is consumed. The growth process continues till the resource runs out and the population faces resource exhaustion. At this stage, as shown in Figure 1(b) the population distribution is broader because of the effect of mutations, and shifted towards lower $\phi$ because of the growth rate function when compared to the initial population. For our simulations we use $r_{d}=0.5$ and $r_{0}=10^{5}$.
It should be noted that although we consider $\phi$ for each cell corresponds to trehalose concentrations produced by the cell eventually in the stationary phase, it also determines its quiescence probability, probability of survival and lag time. Thus, our model implies that mutations accrued in growth process can bias the cell’s performance in other life history stages, although no trehalose is actually produced in the growth phase.
Following experimental observations, upon reaching starvation conditions we assume the stationary phase population differentiates into a heterogeneous population of quiescent and non-quiescent cells in the next simulation phase. This occurs in a single simulation time step. Our model considers the switch to quiescence as a stochastic process, where the probability of each cell in stationary phase going into quiescence is given by
@@
where $\tau_{q}$ is the value of $\phi$ at which probability of quiescence is $50\%$ and $\gamma_{q}$ is a parameter quantifying how quickly the probability rises to 11 as $\phi$ increases. The functional form, as well as the $\tau_{q}$ values we chose allows for probability of quiescence to gradually increase with $\phi$ and saturate at a high value. For each cell we decide the fate of the cell by comparing this probability with a random number drawn from a uniform probability distribution. If the probability given by Equation 2 is higher than the random number the cell goes into a quiescent state. The effect of this probability of quiescence is to impose a lower threshold in $\phi$ on the population, as shown in Figure 1(c).
Entry into stationary phase is followed by exposure to extreme stress, which in the experimental system corresponds to freeze-thaw. This is the third phase in our simulation, consisting of a single simulation time step update. For the purpose of our model we assume that the non-quiescent population entirely perishes upon exposure to extreme stress. Among the quiescent population, cells with higher $\phi$ have a higher probability of survival. We model the probability of survival,
@@
where $\tau_{s}$ is the $\phi$ value at which probability of survival under freeze thaw is $50\%$ and $\gamma_{s}$ is a parameter quantifying how quickly the probability rises to 11 as $\phi$ increases. Again, we choose $\gamma_{s}$ and $\tau_{s}$ such that the probability of survival is close to zero for low values of $\phi$ and then increases relatively quickly in an interval given by $\gamma_{s}$, mimicking a switch-like transition to having enough trehalose production in each cell to lend the cell resistance to freeze-thaw. In general we expect that $\gamma_{s}<\gamma_{q}$ and $\tau_{s}>\tau_{q}$ since the stress is harsher to cells than nutrient scarcity in stationary phase. Thus, both quiescence and the freeze thaw process shift the mean phenotype of the population to higher values since cells with lower phenotype amounts perish. This is shown in Figure 1(d). The functions for growth rate, quiescence probability and survival probability are shown in Figure 2(a).
We model the dilution process before re-introduction into fresh resource of the experimental setup by randomly selecting a fraction $d$ of the population. This can introduce additional stochasticity into the evolution process.
For the final phase of the simulation, when the surviving population is re-introduced into a fresh growth medium, we assume the population remains initially in lag phase, with the lag time of each cell determined by its phenotype. We model cells with higher trehalose concentrations to have a shorter lag time, and resume their growth earlier. We assume the simplest possible functional form, a linear relationship between the lag time and $\phi$,
@@
where $l_{0}$ represents the maximum lag time possible, and $\tau_{l}$ the $\phi$ value at which the lag time reduces to zero. We set $l(\phi)=0$ for all $\phi>\tau_{l}$ for the sake of our simulation, although $\tau_{l}$ is set large enough that $\phi$ does not exceed it in the simulations. Cells also consume a smaller amount of the resource, $r_{l}=0.1$ while they are non-growing in lag phase. Longer lag times thus mean less resource available for growth and the total amount consumed before growth depends on the length of the lag phase. A larger phenotype value thus means a shorter lag phase, and more time available for growth, while simultaneously causing a lower growth rate. The next generation to reach stationary phase, shown in Figure 1(e), is the result of competing effects caused by a phenotype dependent lag time.
@@
Figure 2(b) shows the gradual increase in the mean phenotype of the population as it evolves to stress exposure. When the population starts growth after exit from lag phase with lower values of mean $\phi$, the effect of a decreasing growth rate is not sufficient to stop the increase in mean $\phi$ of the population caused by a shorter lag phase. In the growth phase, cells with higher $\phi$ are able to produce a larger number of offspring and take over the population. This explains the initial increase in the mean phenotype.
@@
Each time the population undergoes a complete round of growth and nutrient exhaustion, we have a resulting heterogeneous population, with cells with higher $\phi$ having a higher probability of entering quiescence. Thus, the distribution of quiescent cells now has a higher mean $\phi$ than the entire stationary population taken together. The cells that are able to survive are a larger proportion of the entire population because of this upward shift in $\phi$. Hence, our simulation includes both a component of phenotype-dependent stochastic switching to quiescence, as well as growth, lag and survival mediated selection for phenotype values. The trade-off between the decrease in growth rate and the positive effect on lag time and survival from increasing trehalose sets the final equilibrium value of the phenotype in the simulations. The effect of a lower growth rate begins to dominate at larger phenotype values, causing the mean phenotype of the population to stagnate.
@@
We see an increase in the phenotype and survival as the evolution process progresses, before settling at an equilibrium value. This is shown in Figure 3. The blue curves represent individual replicate populations, while the red curve represents the average over $100$ replicates. The behavior of populations is stochastic in the early stages of the evolutionary process, due to small numbers of cells that survive freeze-thaw stress, but the equilibrium value they settle to remains conserved across all replicates.
@@
We compare Figure 3 (a) and (b) to the trehalose evolution curve (c) and the survival fraction curve (d) obtained from the experimental evolution of yeast under freeze thaw. Both sets of curves show qualitative agreement. The rapid increase in the survival rate is caused by the large tailed mutational distribution we use which biases evolution towards large jumps. We note also the delay in increase of both the survival curves, which arises, in the case of our simulation, from the sigmoidal dependence of the survival probability upon trehalose concentration.
@@
To determine the effect of model parameters on equilibrium values of the phenotype and survival fraction, we first vary $g_{0}$. This is shown in Figure 4 (a) and (b). Increasing $g_{0}$ increases the number of offspring at each time point, leading to the resource being exhausted faster, thus less time is spent by the population in its growth phase. This causes the effect of selection from growth to decrease, allowing the phenotype to increase to higher values and biasing the population towards maximizing survival. Thus, we see an increase in both the equilibrium value of the phenotype as well as survival fraction as $g_{0}$ increases.
@@
Thus, we find that the amount of time spent in the growth phase, the maximum growth rate, among other parameters can affect the equilibrium value of the phenotype and survival by changing the strength of the trade-off. Note that this hints towards the sensitivity of the final mean phenotype of the population to the experimental setup. We write down the exact form of this dependence in Section III by estimating the long term growth rate of the population.
III Long Term Growth Rate and Comparing Fitness
We calculate the long term growth rate of the evolving population to help us to analytically recover some of the results of our model. Since the life cycle of microbes such as yeast consists of several life history stages, a single trait such as growth rate or survival cannot be taken as a complete indicator of strain fitness. Fitness is a complex combination of these traits, and the long term growth rate can effectively capture this.
Consider a population of cells evolving under growth-stress cycles given by the distribution $n(\phi,t)$. As in the last section, the growth rate, probability of survival under stress and lag time all depend on the phenotype of each cell. Through the growth phase, the number of cells with phenotype $\phi$ at time $(t+1)$ is given by $g(\phi)n(\phi,t)$, where $g(\phi)$ is the growth rate at phenotype $\phi$. If the population is allowed to grow for $T_{g}$ time till the resource runs out, at the end of the growth phase the number of cells at each phenotype is